# OLMES: A Standard for Language Model Evaluations

### v1

@@ Abstract
Progress in AI is often demonstrated by new models claiming improved performance on tasks measuring model capabilities. Evaluating language models in particular is challenging, as small changes to how a model is evaluated on a task can lead to large changes in measured performance. There is no common standard setup, so different models are evaluated on the same tasks in different ways, leading to claims about which models perform best not being reproducible. We propose OLMES, a completely documented, practical, open standard for reproducible LLM evaluations. In developing this standard, we identify and review the varying factors in evaluation practices adopted by the community – such as details of prompt formatting, choice of in-context examples, probability normalizations, and task formulation. In particular, OLMES supports meaningful comparisons between smaller base models that require the unnatural “cloze” formulation of multiple-choice questions against larger models that can utilize the original formulation. OLMES includes well-considered recommendations guided by results from existing literature as well as new experiments investigating open questions. [^1] All prompts, examples, and code used for OLMES are available at https://allenai.org/data/olmes
@@ 1 Introduction
Scientific credibility in AI rests on fair, reproducible comparisons between models. Many current AI models, such as pretrained large language models (LLMs), are generalist models capable of performing downstream tasks they were not specifically trained on (OpenAI,, 2024; Bommasani et al.,, 2022). When evaluating LLMs on such tasks, there are many choices in how the task is presented to the model and how the model outputs are interpreted before scoring (Gao,, 2021; Biderman et al.,, 2024; Liang et al.,, 2023). There is currently no standard way to decide on these choices, and they can have a huge impact on model performance, with some recent papers claiming as much as an 80% difference in accuracy on a given task just from varying formatting and in-context examples (Sclar et al.,, 2023). These choices are rarely reported with enough information to actually reproduce, so when a team of ML practitioners releases a new model it is often impossible for them to compare against previously-reported results.
Efforts like the Holistic Evaluation of Language Models (HELM) benchmark (Liang et al.,, 2023) and the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023) strive towards standardizing LM evaluations. While the same setup is used to evaluate many models ensuring consistency and reproducibility, the rationale behind prompt formatting, use of in-content examples, normalization techniques, and task formulation are not always clearly documented and thus not consistently followed by other researchers in subsequent work (Touvron et al., 2023b, ; Biderman et al.,, 2023; Jiang et al.,, 2023; Groeneveld et al.,, 2024; AI@Meta,, 2024).
We highlight two problems in the field today. (1) Releasing a new model and comparing against previously reported results is flawed unless the previous work explicitly described their full evaluation setup, and then that setup is followed in the new work. Unfortunately, such results are not comparable to previous work that used a different format. Importantly, the key goal of comparing models on the same downstream task is to measure their relative strengths. This is a real problem, as we show in Table 1, which illustrates how several models’ published performance on the ARC-Challenge (Clark et al.,, 2018) and OpenbookQA (Mihaylov et al.,, 2018) tasks varies in the literature. For instance, looking at Ref1 for ARC-Challenge, we would conclude that Llama2-13B and Llama3-8B are performing similarly, but Ref6 reveals there is likely a gap of over 10% between them. For OpenbookQA, on Falcon-7B, there is a discrepancy of more than 25% between Ref2 and Ref8, which only makes sense if one considers the very different task formulations. (2) Some papers are likely reporting over-estimates of performance by optimizing the prompt format for their model, while others are not. In fact, we see evidence that this might be a community-wide problem in a recent effort from HELM (Mai and Liang,, 2024). When evaluating a suite of models using a single format and comparing against each model’s self-reported result, they find significant differences in performance. In particular, their findings suggest that models with higher self-reported performance were more likely to have lower scores in the reproduction, suggesting some over-estimating of model performance (see Figure 4 in Appendix).
Table 1: Scores reported in different references for LLM performances on ARC-Challenge and OpenbookQA. Scores indicated with are using multiple-choice formulation (MCF) rather than “cloze” formulation (CF) (see Section 2.1 for definitions). Entries with “?” denote either undocumented or mixed approaches across models. Different references use different evaluation setups, some of which are not fully specified, so conclusions about which models perform best are not reproducible.
ARC-Challenge Evaluations: OpenbookQA Evaluations:
@@
MPT-7B 47.7 42.6 46.5 45.7 51.4 48.6 52.4
RPJ-INCITE-7B 46.3 42.8 45.3 49.4 49.0
Falcon-7B 47.9 42.4 44.5 47.5 49.7 51.6 44.6 53.0 $26.0^{\dagger}$ 55.2
Num shots 25 0 0 0 0 25 5 0 0 0 0 5 5
Curated shots No No Yes No Yes
Formulation RC RC RC? RC RC MC MC/RC RC RC RC RC MC MC/RC
Ref Reference citation Ref Reference citation
Ref1 HF Open LLM Leaderboard (Beeching et al.,, 2023) Ref5 OLMo paper (Groeneveld et al.,, 2024)
Ref2 Llama2 paper (Touvron et al., 2023a, ) Ref6 Llama3 model card (AI@Meta,, 2024)
Ref3 Mistral 7B (Jiang et al.,, 2023) Ref7 Gemma paper (Gemma Team et al.,, 2024)
Ref4 Falcon paper (Almazrouei et al.,, 2023) Ref8 HELM Lite Leaderboard (Liang et al.,, 2023)
To address these problems, we present OLMES (Open Language Model Evaluation Standard), a standard to improve the transparency and reproducibility of language model evaluation from a practical point of view, removing ambiguity in how a final performance metric is obtained when evaluating a model on a dataset. OLMES can be applied to evaluation during the model development process, leaderboards, papers, providing justified recommendations on data sampling, how to format instances, the choice of in-context examples, probability normalization, and task formulation.
@@
  • Documented: Each decision in the standard is documented with justifications by applying principles from existing studies and performing experiments to investigate open questions.
@@
Since OLMES is a documented, practical, open evaluation standard, it is straightforward to adopt in publicly available, well-maintained evaluation code bases like the Eleuther LM Evaluation Harness (Gao et al.,, 2023; Biderman et al.,, 2024) and HELM (Liang et al.,, 2023). When used by model developers and other researchers, OLMES will help unify evaluation practices in the field. All prompts, examples, and code used for OLMES can be found at https://allenai.org/data/olmes.
@@ 2.1 Multiple-choice QA and LLM evaluation
Multiple-choice question answering (MCQA) tasks present a compelling way of evaluating models and humans alike, due to the ease of scoring (whether the correct answer is chosen out of the given options) and the allowed flexibility in the domain and complexity of the questions. Such tasks form a large, essential part of LM evaluations and are be the focus in OLMES. There are generally two ways to formulate these tasks.
@@
presenting answer choices indicated by labels and scoring prediction of answer labels, just like how MCQA is posed to humans. Here is an example of MCQA from ARC-Easy (Clark et al.,, 2018), a dataset of real grade-school level science questions:
Question: Earth’s core is primarily composed of which of the following materials? (A) basalt (B) iron (C) magma (D) quartz
Answer: (B)
@@
scoring each answer choice separately using LM token probabilities. The MCF format is not compatible with the pure language modeling nature of generating the next token. Therefore, the CF format was introduced by when evaluating the GPT-3 model (Brown et al.,, 2020). They found that it was possible to elicit much better performance using a “cloze” completion version of the task, where the model is shown a prompt like:
@@
Answer: <answer>,
where each answer choice is separately substituted in for <answer>.
Then the LM probability of the answer choice tokens are used to rank the choices and predict an answer. This formulation has natural limitations, such as not being able to properly address cases where one answer choice is “none of the above”.
@@ 2.2 Targeted tasks
Table 2: OLMES details on tasks, with our standardized choices of dataset split, number of instances to use (along with total number if sampling was used), and which CF normalization scheme to use (see Section 3.3). Column #C shows the number of answer choices (ARC-Challenge and ARC-Easy have a few instances with 3 or 5 answer choices).
@@
ARC-Challenge (ARC_C) Test $4^{\dagger}$ 1172 pmi (Clark et al.,, 2018)
ARC-Easy (ARC_C) Test $4^{\dagger}$ 1000 (2376) char (Clark et al.,, 2018)
BoolQ Val 2 1000 (3270) none (Clark et al.,, 2019)
CommonsenseQA (CSQA) Val 5 1221 pmi (Talmor et al.,, 2019)
HellaSwag (HSwag) Val 4 1000 (10042) char (Zellers et al.,, 2019)
MMLU Test 4 14042 char (Hendrycks et al.,, 2021)
OpenbookQA (OBQA) Test 4 500 pmi (Mihaylov et al.,, 2018)
PIQA Val 2 1000 (1838) char (Bisk et al.,, 2020)
Social IQa (SIQA) Val 3 1000 (1954) char (Sap et al.,, 2019)
WinoGrande (WinoG) Val 2 1267 none (Sakaguchi et al.,, 2020)
We select and implement standards for 10 popular benchmark MCQA tasks, see Table 2 for the list. The list covers tasks that are frequently used in the community’s evaluation practices, such as the Hugging Face Open LLM Leaderboard (Beeching et al.,, 2023), Llama papers (Touvron et al., 2023a, ; Touvron et al., 2023b, ; AI@Meta,, 2024), HELM (Liang et al.,, 2023), and the OLMo evaluation suite (Groeneveld et al.,, 2024). This selection includes questions on science, various types of commonsense, factual knowledge, and covers a range of topics (MMLU alone covers 57 subjects), of varying difficulty.
@@ 2.3 Selection of models
We develop OLMES based on a selection of 15 diverse, openly available pretrained LLMs, covering a range of sizes from 1B to 70B – Pythia-1B, Pythia-6.7B (Biderman et al.,, 2023), OLMo-1B, OLMo-7B, OLMo-1.7-7B (Groeneveld et al.,, 2024), TinyLlama-1.1B (Zhang et al.,, 2024), StableLM2-1.6B (Bellagente et al.,, 2024), RPJ-INCITE-7B (Together Computer,, 2023), MPT-7b (MosaicML,, 2023), Falcon-7B (Almazrouei et al.,, 2023), Llama2-7B, Llama2-13B (Touvron et al., 2023b, ), Mistral-7B-v0.1 (Jiang et al.,, 2023), Llama3-8B, Llama3-70B (AI@Meta,, 2024). This reflects our goal of providing an evaluation standard that suits a range of model capabilities, with the flexibility to apply the same methodology during model development as well as when comparing final powerful models.
@@ 3 Standardizing variations in evaluation
To evaluate a model on a dataset, there are a variety of decisions that have to be made to get a final accuracy metric of that model on that dataset. These include:
@@
  • Other implementation choices impacting results (Appendix A)
Below we enumerate key variations in these steps, and justify the choices made in OLMES (listed in Table 2) to standardize these steps, leaving some of the details for the Appendix.
@@ 3.1 How to format dataset instances?
Each MCQA dataset includes a set of fields used to specify an instance, such as question, answer choices, and perhaps a context for the question. When formatting an instance as a prompt to an LLM, many different choices have been made in the literature. This includes simple choices like "Question:" vs "Q:" as question prefix (varying even within a paper, e.g., Brown et al., (2020)), or formatting the answer labels (e.g., "A." (Touvron et al., 2023a, ), "(A)" (Nori et al.,, 2023), "<mc>A</mc>" (Anthropic,, 2024), etc). There is also a choice of whether or not to provide a general instruction (e.g., common for MMLU (Hendrycks et al.,, 2021), sometimes done for OpenbookQA (Almazrouei et al.,, 2023)).
@@
Sampling. Following existing LM evaluation standardization efforts (Liang et al.,, 2023; Beeching et al.,, 2023), OLMES uses the test split of a dataset if the labels are publicly available, otherwise the validation split. If the dataset has more than 1500 instances, we sample 1000 instances to evaluate on, [^2] Sampling uses a specific random seed in Python: Random(1234).sample(all_instances, 1000) similar to HELM (Liang et al.,, 2023) which caps evaluation instances at 1000.333https://crfm-helm.readthedocs.io/en/latest/reproducing_leaderboards/ Note that the potential extra statistical signal from more instances would generally be dominated by other sources of score variations, like prompt formatting, so this is a practical consideration to avoid unnecessary computation resources. See Table 2 for details on splits and sampling used in OLMES.

### v2

@@ Abstract
Progress in AI is often demonstrated by new models claiming improved performance on tasks measuring model capabilities. Evaluating language models can be particularly challenging, as choices of how a model is evaluated on a task can lead to large changes in measured performance. There is no common standard setup, so different models are evaluated on the same tasks in different ways, leading to claims about which models perform best not being reproducible. We propose OLMES, a completely documented, practical, open standard for reproducible LLM evaluations. In developing this standard, we identify and review the varying factors in evaluation practices adopted by the community – such as details of prompt formatting, choice of in-context examples, probability normalizations, and task formulation. In particular, OLMES supports meaningful comparisons between smaller base models that require the unnatural “cloze” formulation of multiple-choice questions against larger models that can utilize the original formulation. OLMES includes well-considered, documented recommendations guided by results from existing literature as well as new experiments resolving open questions. [^1] All prompts, examples, and code used for OLMES are available at https://github.com/allenai/olmes.
OLMES: A Standard for Language Model Evaluations
Yuling Gu α  Oyvind Tafjord α  Bailey Kuehl α  Dany Haddad α Jesse Dodge α  Hannaneh Hajishirzi αβ αAllen Institute for Artificial Intelligence  βUniversity of Washington {yulingg, oyvindt}@allenai.org
@@ 1 Introduction
Scientific credibility in AI rests on reproducible and well-considered comparisons between models. Many current AI models, such as pretrained large language models (LLMs), are generalist models capable of performing downstream tasks they were not specifically trained on (Brown et al., 2020; Bommasani et al., 2022). When evaluating LLMs on such tasks, there are many choices in how the task is presented to the model and how the model outputs are interpreted before scoring (Gao, 2021; Biderman et al., 2024; Liang et al., 2023). There is currently no standard way to decide on these choices, and they can have significant impact on model performance, with some recent papers claiming as much as an 80% difference in accuracy on a given task just from varying formatting and in-context examples (Sclar et al., 2023).
@@
MPT-7B 47.7 42.6 46.5 45.7
RPJ-Incite-7B 46.3 42.8 45.3
Falcon-7B 47.9 42.4 44.5 47.5 49.7
Mistral-7B 60.0 55.5 54.9 $78.6^{\dagger}$
Llama2-7B 53.1 45.9 43.2 45.9 48.5 $53.7^{\dagger}$ 54.2
Llama2-13B 59.4 49.4 48.8 49.4 $67.6^{\dagger}$ $67.3^{\dagger}$
Llama3-8B 60.2 $78.6^{\dagger}$ $79.3^{\dagger}$
Num shots 25 0 0 0 0 25 5
Curated shots No No Yes
Formulation CF CF CF? CF CF MCF MCF/CF
Normalization char char ? char? pmi none none/pmi
Ref Reference citation
Ref1 HF Open LLM Leaderboard (Beeching et al., 2023)
Ref2 Llama2 paper (Touvron et al., 2023a)
Ref3 Mistral 7B (Jiang et al., 2023)
Ref4 Falcon paper (Almazrouei et al., 2023)
Ref5 OLMo paper (Groeneveld et al., 2024)
Ref6 Llama3 model card (AI@Meta, 2024)
Table 1: Scores reported in different references for LLM performances on ARC-Challenge. Scores indicated with are using multiple-choice formulation (MCF) rather than “cloze” formulation (CF) (see Section 2.1 for definitions). Entries with “?” denote either undocumented or mixed approaches across models. Different references use different evaluation setups, some of which are not fully specified, so conclusions about performances and relative strengths of models are not reproducible.
These choices in evaluation setups are often not reported with enough details to reproduce, so when a team of ML practitioners releases a new model it is often impossible for them to directly compare against previously-reported results by others. Efforts like the Holistic Evaluation of Language Models (HELM) benchmark (Liang et al., 2023) and the Hugging Face Open LLM Leaderboard (Beeching et al., 2023) tackle the issue of reproducibility by striving towards standardizing LM evaluations. While the same setup is used to evaluate many models ensuring consistency and reproducibility, the rationale behind prompt formatting, use of in-content examples, normalization techniques, and task formulation are not always clearly documented and thus not consistently followed by other researchers in subsequent work (Touvron et al., 2023b; Biderman et al., 2023; Jiang et al., 2023; Groeneveld et al., 2024; AI@Meta, 2024).
We highlight two problems in the field today. (1) Releasing a new model and comparing it against previously reported results is flawed unless the previous work explicitly described their full evaluation setup, and then that setup is followed in the new work. Currently, different references (like leaderboards, papers) use different (and sometimes under-specified) evaluation setups, leading to different results and conclusions. For a particular dataset, evaluating a specified language model, different references can tell you very different stories. We illustrate this phenomenon in Table 1, which shows how several models’ published performance on the ARC-Challenge (Clark et al., 2018) task can vary in the literature. For instance, looking at Ref1, we would conclude that Llama2-13B and Llama3-8B are performing similarly, but Ref6 reveals there is likely a gap of over 10% between them. (2) Despite current efforts to standardize model evaluation (e.g., HF Open LLM Leaderboard, HELM), the choices made are not justified and most model creators do not use these setups for their evaluations. We also see evidence of this in Table 1, showing a variety of setups being used for the same task, differing in choices such as number of shots, source of in-context examples, task formulation, and probability normalization. While these different choices are made in implementing the evaluations, to date, there is no documented standard studying and/or justifying if one choice is better than another, leading to the lack of a set of justified choices that the community can adopt. Mai and Liang (2024) also demonstrates how this might be a community-wide problem in a recent effort (see Figure 4 in Appendix).
To address these problems, we present OLMES (Open Language Model Evaluation Standard), a standard to improve the transparency and reproducibility of language model evaluation from a practical point of view, removing ambiguity in how a final performance metric is obtained when evaluating a model on a dataset. OLMES can be applied to evaluation during the model development process, and in published leaderboards and papers. OLMES provides justified recommendations on all aspects of task setups, such as data sampling, how to format instances, the choice of in-context examples, probability normalization, and task formulation.
@@
  • Documented: Each decision in the standard is documented with justifications by applying principles from existing studies and performing experiments to resolve open questions.
@@
Since OLMES is a documented, practical, open evaluation standard, it is straightforward to adopt in publicly available, well-maintained evaluation code bases like the Eleuther LM Evaluation Harness (Gao et al., 2023; Biderman et al., 2024) and HELM (Liang et al., 2023). When used by model developers and other researchers, OLMES will help unify evaluation practices in the field. We believe this work is the first of its kind to unify practices for evaluating base models throughout the full development cycle, from small to large models as well as early to late training stages. All prompts, examples, and code used for OLMES can be found at https://github.com/allenai/olmes.
@@ 2.1 Multiple-choice QA and LLM evaluation
Multiple-choice question answering (MCQA) tasks present a compelling way of evaluating models and humans alike, due to the ease of scoring (whether the correct answer is chosen out of the given options) and the allowed flexibility in the domain and complexity of the questions. One motivation for multiple-choice tasks is that early in training, and for smaller base models before instruction-tuning, other tasks (generative tasks, math reasoning, coding, etc) tend to provide less useful signals. Multiple-choice tasks are the most common type of benchmarks for evaluating base LLMs (Beeching et al., 2023; Touvron et al., 2023a; Jiang et al., 2023; Groeneveld et al., 2024; AI@Meta, 2024), where the evaluation seems straightforward (did the model predict the right answer?), but in practice, a statement like “model X scores Y on ARC-Challenge” is generally uninterpretable (with unspecified details and cannot be meaningfully compared across references, see Table 1) without a clear evaluation standard like OLMES.
We specifically focus on evaluation using these tasks to provide useful guidance during and after base model training, giving important insights into the potential of such models before committing to further tuning (e.g., instruction-tuning). Such tasks form a large, essential part of LLM evaluations and are the focus of OLMES. There are generally two ways to formulate these tasks.
@@
presenting answer choices indicated by labels and scoring prediction of answer labels, just like how MCQA is posed to humans. Here is an example of MCQA from ARC-Easy (Clark et al., 2018), a dataset of real grade-school level science questions:
   A. basalt
   B. iron
   C. magma
   D. quartz
Answer: B
@@
scoring each answer choice separately using LLM token probabilities. The MCF format is not natural for the pure language modeling task of generating the next token. Therefore, the CF format was introduced when evaluating the GPT-3 model (Brown et al., 2020). They found that it was possible to elicit much better performance using a “cloze” completion version of the task, where the model is shown a prompt like:
@@
Answer: <answer>
Each answer choice is separately substituted in for <answer>. Then the LLM probability of the answer choice tokens are used to rank the choices and predict an answer. This formulation has ambiguities in how to normalize the probability, as well as absolute limitations, such as not being able to properly address cases where one answer choice is “none of the above” or similar.
@@ 2.2 Targeted tasks
ARC-Challenge (ARC_C) Test $4^{\dagger}$ 1172 pmi (Clark et al., 2018)
ARC-Easy (ARC_E) Test $4^{\dagger}$ 1000 (2376) char (Clark et al., 2018)
BoolQ Val 2 1000 (3270) none (Clark et al., 2019)
CommonsenseQA (CSQA) Val 5 1221 pmi (Talmor et al., 2019)
HellaSwag (HSwag) Val 4 1000 (10042) char (Zellers et al., 2019)
MMLU Test 4 14042 char (Hendrycks et al., 2021)
OpenbookQA (OBQA) Test 4 500 pmi (Mihaylov et al., 2018)
PIQA Val 2 1000 (1838) char (Bisk et al., 2020)
Social IQa (SIQA) Val 3 1000 (1954) char (Sap et al., 2019)
WinoGrande (WinoG) Val 2 1267 none (Sakaguchi et al., 2020)
Table 2: OLMES details on tasks, with our standardized choices of dataset split, number of instances to use (along with total number if sampling was used), and which CF normalization scheme to use (see Section 3.3). Column #C shows the number of answer choices (ARC-Challenge and ARC-Easy have a few instances with 3 or 5 answer choices). See Section 3 for details on instance formatting, choice of in-context examples and task formulation.
We select and implement standards for 10 popular benchmark MCQA tasks, see Table 2 for the list. The list covers tasks that are frequently used in the community’s evaluation practices, such as the Hugging Face Open LLM Leaderboard (Beeching et al., 2023), Llama papers (Touvron et al., 2023a, b; AI@Meta, 2024), HELM (Liang et al., 2023), and the OLMo evaluation suite (Groeneveld et al., 2024). This selection includes questions on science, various types of commonsense, factual knowledge, and covers a range of topics (MMLU alone covers 57 subjects), of varying difficulty.
@@ 2.3 Selection of models
We develop OLMES based on a selection of 15 diverse, openly available pretrained LLMs, focusing on base (not instruction-tuned) models, covering a range of sizes from 1B to 70B – Pythia-1B, Pythia-6.7B (Biderman et al., 2023), OLMo-1B, OLMo-7B, OLMo-7B-0424 (Groeneveld et al., 2024), TinyLlama-1.1B (Zhang et al., 2024), StableLM2-1.6B (Bellagente et al., 2024), RPJ-INCITE-7B (Together Computer, 2023), MPT-7b (MosaicML, 2023), Falcon-7B (Almazrouei et al., 2023), Llama2-7B, Llama2-13B (Touvron et al., 2023b), Mistral-7B-v0.1 (Jiang et al., 2023), Llama3-8B, Llama3-70B (AI@Meta, 2024). This reflects our goal of providing an evaluation standard that suits a range of model capabilities, with the flexibility to apply the same methodology during model development as well as when comparing final powerful base models.
Assessing base models of different strengths is important during the training of models and before it is used for further tuning (e.g., instruction-tuning). This is critical for the community when picking between alternate base models for further training or tuning for their application. There is limited established protocol in the community – evaluation during training is often left underspecified and understudied, and when evaluating final base models, researchers across the field use different evaluation setups, leading to different results and conclusions (Tables 1 and 14). We hope this work will empower the community towards more unified practices in benchmarking base models so that further progress can be made on a stronger foundation based on careful evaluation.
@@ 3 Standardizing variations in evaluation
To evaluate a model on a dataset, there are a variety of decisions that have to be made to get a final score of that model on that dataset. These include:
@@
  • Other implementation choices impacting results (Section 3.5)
Below we enumerate key variations in these steps, and justify the choices made in OLMES (some of which are summarized in Table 2) to standardize these steps, leaving some of the details for the Appendix.
@@ 3.1 How to format dataset instances?
Each MCQA dataset includes a set of fields used to specify an instance, such as question, answer choices, and perhaps a context for the question. When formatting an instance as a prompt to an LLM, many different choices have been made in the literature. This includes simple choices like "Question:" vs "Q:" as question prefix (varying even within a paper, e.g., Brown et al. (2020)), or formatting the answer labels (e.g., "A." (Touvron et al., 2023a), "(A)" (Nori et al., 2023), "<mc>A</mc>" (Anthropic, 2024), etc). There is also a choice of whether or not to provide a general instruction, e.g., common for MMLU (Hendrycks et al., 2021), sometimes done for OpenbookQA (Almazrouei et al., 2023).
@@
Sampling. Following existing LLM evaluation standardization efforts (Liang et al., 2023; Beeching et al., 2023), OLMES uses the test split of a dataset if the labels are publicly available, otherwise the validation split. If the dataset has more than 1500 instances, we sample 1000 instances to evaluate, [^2] Sampling uses a specific random seed in Python: Random(1234).sample(all_instances, 1000) similar to HELM (Liang et al., 2023) which caps evaluation instances at 1000.333https://crfm-helm.readthedocs.io/en/latest/reproducing_leaderboards/ Note that the potential extra statistical signal from more instances would generally be dominated by other sources of score variations, like prompt formatting, so this is a practical consideration to avoid unnecessary computation resources. See Table 2 for details on splits and sampling used in OLMES.